OpenStudy (anonymous):

The table below shows the ticket rates for whale watching trips offered by Ocean Tours: Age Ticket Price Under 3 years free 3 to 12 years $36 Over 12 years $48 On a certain day, the company took 92 people on whale watching trips. There were 40 children aged 12 and under, of which some children were under 3 years. If x represents the number of children under 3 years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day? (4 points)

OpenStudy (anonymous):

Possible Answers: (92 - x)36 + (92 - 40) ⋅ 48 = C (40 + x)36 + (92 - (40 + x)) ⋅ 48 = C (92 + x)36 + (92 - (40 + x)) ⋅ 48 = C (40 - x)36 + (92 - 40) ⋅ 48 = C

OpenStudy (anonymous):

(40-x)(36) + (92 - 40) * 48 = C There were 92 people total, 40 of which were under 12, leaving the (92-40) people over 12. Each ticket for them cost 48 dollars, so the revenue for people over 12 is (92-40)(48) There's an unknown number of people that's under 3. The kids that are under 3 do not generate revenue, so they're not a part of the cost. Essentially, they are x * 0 = 0 anyways

OpenStudy (anonymous):

Okay :) Thanks!